you have two kids and one says to the other if you took two years off my age and gave it to you, you would be double my age and if you took one more year you would be triple my age. How old are the kids?
Found 2 solutions by solver91311, RAY100:
Answer by solver91311(24713)



Let x represent the age of the one doing the talking. If you take two years off of that age you have x - 2. Let y represent the age of the the other one. If you add two years to that age you have y + 2.

After the one doing the talking has given away the two years:



Likewise, if you take three years off of the talker's age you have x - 3. If you add two years to the other one's age you have y + 3. And:



Put the equations in standard form:





Multiply the first equation by -1 and then add the equations:







The one doing the talking is 6, so:





And the other one is 6 as well. Hmmm, twins, and mighty clever for 6 year olds.

Check:

6 minus 2 is 4, 6 plus 2 is 8, 8 is twice 4.

6 minus 3 is 3, 6 plus 3 is 9, 9 is three times 3.

John


Answer by RAY100(1637)

let x, y =age of kids
1)x+2=2(y-2)=2y-4,,,,x-2y=-6
2)x+3=3(y-3)=3y-9,,,,x-3y=-12
subt(2)-(1)
y=6
subst into (1)
x-2(6)=-6
x=6
check(1) 6+2=2(6-2),,8=8 ,,,ok
check(2) 6+3=3(6-3),,9=9,,,ok
therefore both kids are 6